# Smooth Tests for Normality in ANOVA

### v3

@@ Appendix A Appendix
Under $H_{0}$, by CLT, $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})\to_{d}\mathcal{N}(0,2\sigma^{4})$ as $N\to\infty$, then $\hat{\sigma}^{2}-\sigma^{2}=O_{p}(N^{-1/2})$.
Now we deal with $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$. By second order Taylor expansion of $\pi_{k}(\hat{Z}_{ij})$ with respect to $\hat{Z}_{ij}$ at $Z_{ij}$, we obtain
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\left(\hat{Z}_{ij}-Z_{ij}\right)$
$+\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\ddot{\pi}_{k}\left(\tilde{Z}_{ij}\right)\left(\hat{Z}_{ij}-Z_{ij}\right)^{2},$ (A.3)
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)+\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\dot{\phi}\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{2},$ (A.4)
where $\tilde{\varepsilon}_{ij}/\tilde{\sigma}$ an intermediate point between $^{\hat{\varepsilon}_{ij}}/{\hat{\sigma}}$ and $\varepsilon_{ij}/\sigma$, whose exact value may vary across different scenarios.
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}$
$=$ $-\frac{\hat{\mu}-\mu}{\hat{\sigma}}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)-\frac{\hat{\sigma}^{2}-\sigma^{2}}{\hat{\sigma}(\hat{\sigma}+\sigma)}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$
$=$ $-\frac{\hat{\mu}-\mu}{\sigma}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]-\frac{\hat{\sigma}^{2}-\sigma^{2}}{2\sigma^{2}}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]$
$+o_{p}\left(\hat{\mu}-\mu\right)+o_{p}\left(\hat{\sigma}^{2}-\sigma^{2}\right),$ (A.5)
where the last step follows due to the law of large numbers (LLN) for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})=\mathrm{E}[\dot{\pi}_{k}(Z)\phi(e)]+o_{p}(1)$ and $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}=\mathrm{E}[\dot{\pi}_{k}(Z)\phi(e)e]+o_{p}(1)$ as $N\to\infty$, as well as the consistency of $\hat{\sigma}$ to $\sigma$ from the arguments above. Under $H_{0}$, through integration by parts, we have
$\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=\int_{0}^{1}\pi_{k}(z)\Phi^{-1}(z)\mathrm{d}z=c_{1k},$
@@
$\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=\int_{0}^{1}\pi_{k}(z)\left[\left(\Phi^{-1}(z)\right)^{2}-1\right]\mathrm{d}z=\int_{0}^{1}\pi_{k}(z)\left(\Phi^{-1}(z)\right)^{2}\mathrm{d}z=c_{2k},$
@@
$R_{N1}\equiv\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\ddot{\pi}_{k}\left(\tilde{Z}_{ij}\right)\left(\hat{Z}_{ij}-Z_{ij}\right)^{2}=o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.7)
@@
$R_{N2}\equiv\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\dot{\phi}\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.8)
@@
$|R_{N1}|$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left[\Phi\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}\right)-\Phi\left(\frac{{\varepsilon}_{ij}}{\sigma}\right)\right]^{2}$
$=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\phi^{2}\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
$\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
$=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}.$
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}$
$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}$
$\asymp$ $\left(\hat{\mu}-\mu\right)^{2}+(\hat{\sigma}^{2}-\sigma^{2})^{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}$
@@
$|R_{N2}|$ $=\left|\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\left[-\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\phi\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\right]\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}\right|$
$\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left|\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right|\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
$\leq\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}^{2}}\right)^{1/2}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{4}\right)^{1/2},$
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{4}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{4}$
$\lesssim$ $\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{4}\right]$
$\asymp$ $\left(\hat{\mu}-\mu\right)^{4}+\left(\hat{\sigma}^{2}-\sigma^{2}\right)^{4}\left(\mathrm{E}\left[e^{4}\right]+o_{p}(1)\right)=O_{p}\left(\frac{1}{N^{2}}\right),$
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}^{2}}$ $\leq\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{{\varepsilon}_{ij}}{{\sigma}}\right)^{2}$
@@
Proof of Theorem 2.
The proof of Theorem 2 is similar to that of Theorem 1. Under $H_{1}$, (A.1) still holds, and $\sqrt{N}(\hat{\mu}-\mu)\to_{d}\mathcal{N}(0,\sigma^{2})$ since $e_{ij}\sim F\neq\Phi$. Analogously, (A.2) still holds, and $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})\to_{d}\sigma^{2}\mathcal{N}(0,\mathrm{E}[e^{4}]-1)$.

### v4

@@ A.1 Proof of Theorem 1
Under $H_{0}$, by CLT, $\sqrt{N}(\widehat{\sigma}^{2}-\sigma^{2})\overset{d}{\to}\mathcal{N}(0,2\sigma^{4})$ as $N\to\infty$, then $\widehat{\sigma}^{2}-\sigma^{2}=O_{p}(N^{-1/2})$.
Now we deal with $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$. By second order Taylor expansion of $\pi_{k}(\widehat{Z}_{ij})$ with respect to $\widehat{Z}_{ij}$ at $Z_{ij}$, we obtain
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\left(\widehat{Z}_{ij}-Z_{ij}\right)$
$+\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\ddot{\pi}_{k}\left(\widetilde{Z}_{ij}\right)\left(\widehat{Z}_{ij}-Z_{ij}\right)^{2},$ (A.3)
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)+\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\dot{\phi}\left(\frac{\widetilde{\varepsilon}_{ij}}{\widetilde{\sigma}}\right)\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{2},$ (A.4)
where $\widetilde{\varepsilon}_{ij}/\widetilde{\sigma}$ an intermediate point between $^{\widehat{\varepsilon}_{ij}}/{\widehat{\sigma}}$ and $\varepsilon_{ij}/\sigma$, whose exact value may vary across different scenarios.
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}$
$=$ $-\frac{\widehat{\mu}-\mu}{\widehat{\sigma}}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)-\frac{\widehat{\sigma}^{2}-\sigma^{2}}{\widehat{\sigma}(\widehat{\sigma}+\sigma)}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$
$=$ $-\frac{\widehat{\mu}-\mu}{\sigma}\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]-\frac{\widehat{\sigma}^{2}-\sigma^{2}}{2\sigma^{2}}\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]$
$+o_{p}\left(\widehat{\mu}-\mu\right)+o_{p}\left(\widehat{\sigma}^{2}-\sigma^{2}\right),$ (A.5)
where the last step follows due to the law of large numbers (LLN) for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})=\mathbb{E}[\dot{\pi}_{k}(Z)\phi(e)]+o_{p}(1)$ and $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}=\mathbb{E}[\dot{\pi}_{k}(Z)\phi(e)e]+o_{p}(1)$ as $N\to\infty$, as well as the consistency of $\widehat{\sigma}$ to $\sigma$ from the arguments above. Under $H_{0}$, through integration by parts, we have
$\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=\int_{0}^{1}\pi_{k}(z)\Phi^{-1}(z)\mathrm{d}z=c_{1k},$
@@
$\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=\int_{0}^{1}\pi_{k}(z)\left[\left(\Phi^{-1}(z)\right)^{2}-1\right]\mathrm{d}z=\int_{0}^{1}\pi_{k}(z)\left(\Phi^{-1}(z)\right)^{2}\mathrm{d}z=c_{2k},$
@@
$R_{N1}\equiv\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\ddot{\pi}_{k}\left(\widetilde{Z}_{ij}\right)\left(\widehat{Z}_{ij}-Z_{ij}\right)^{2}=o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.7)
@@
$R_{N2}\equiv\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\dot{\phi}\left(\frac{\widetilde{\varepsilon}_{ij}}{\widetilde{\sigma}}\right)\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.8)
@@
$|R_{N1}|$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left[\Phi\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}\right)-\Phi\left(\frac{{\varepsilon}_{ij}}{\sigma}\right)\right]^{2}$
$=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\phi^{2}\left(\frac{\widetilde{\varepsilon}_{ij}}{\widetilde{\sigma}}\right)\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
$\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
$=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}\right)^{2}.$
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}\right)^{2}$
$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}\right)^{2}$
$\asymp$ $\left(\widehat{\mu}-\mu\right)^{2}+(\widehat{\sigma}^{2}-\sigma^{2})^{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}$
@@
$|R_{N2}|$ $=\left|\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\left[-\frac{\widetilde{\varepsilon}_{ij}}{\widetilde{\sigma}}\phi\left(\frac{\widetilde{\varepsilon}_{ij}}{\widetilde{\sigma}}\right)\right]\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}\right|$
$\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left|\frac{\widetilde{\varepsilon}_{ij}}{\widetilde{\sigma}}\right|\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
$\leq\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\widetilde{\varepsilon}_{ij}^{2}}{\widetilde{\sigma}^{2}}\right)^{1/2}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{4}\right)^{1/2},$
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{4}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}\right)^{4}$
$\lesssim$ $\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}\right)^{4}\right]$
$\asymp$ $\left(\widehat{\mu}-\mu\right)^{4}+\left(\widehat{\sigma}^{2}-\sigma^{2}\right)^{4}\left(\mathbb{E}\left[e^{4}\right]+o_{p}(1)\right)=O_{p}\left(\frac{1}{N^{2}}\right),$
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\widetilde{\varepsilon}_{ij}^{2}}{\widetilde{\sigma}^{2}}$ $\leq\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{{\varepsilon}_{ij}}{{\sigma}}\right)^{2}$
@@
A.2 Proof of Theorem 2
The proof of Theorem 2 is similar to that of Theorem 1. Under $H_{1}$, (A.1) still holds, and $\sqrt{N}(\widehat{\mu}-\mu)\overset{d}{\to}\mathcal{N}(0,\sigma^{2})$ since $e_{ij}\sim F\neq\Phi$. Analogously, (A.2) still holds, and $\sqrt{N}(\widehat{\sigma}^{2}-\sigma^{2})\overset{d}{\to}\sigma^{2}\mathcal{N}(0,\mathbb{E}[e^{4}]-1)$.